# Representation Bending for Large Language Model Safety

### v2

@@ 1 Introduction
Figure 1: RepBend bends the representations of the model to be far from the unsafe representations and close to the safe representations, while maintaining its general capabilities.
@@ 3 Representation Bending
The core idea of representation bending is to “bend” the representation of the model in a way to make the model closer to safer states – latent knowledge encoded in the activation space – while making it further from harmful states. In other words, remap model representations related to harmful states, redirecting them towards incoherent or refusal representations (Figure 1).
@@
RepBend is shown in Algorithm 1. RepBend tries to change the representations of model $M$ by minimizing a loss function over a fixed number of steps during fine-tuning. At a high level, RepBend bends the representations of model to be far from unsafe representations and close to safe representations, while maintaining its general capabilities. For improved performance, RepBend updates the model parameters with LoRA Hu et al. (2021). Model with LoRA parameters is referred to as (and can be seen as an independent) model $M^{\prime}$. Since we only update the LoRA parameters, $M^{\prime}$ at the end of the algorithm would be the desired safe model.
@@
Line 12 shows four terms. in the loss function. 1. First loss term keeps the model’s representation close to the safe representations (small L2 difference); this can be seen as “retain loss.” 2. Second loss term is the opposite of the first one: it pushes the representations of the model to be far away from unsafe representations (large L2 difference); this can be seen as “forget loss” somehow. 3. The cosine similarity loss term stabilizes the model responses under the unsafe queries toward a single “refusal-like” direction (e.g., by terms like “I am sorry I cannot help” or “As an AI assistant, I am unable to …”), helping the model consistently reject them rather than producing random or inconsistent outputs. Recall that cosine similarity is high when inputs are similar, and we want similarity, hence the negative sign for the loss. 4. Finally, the KL divergence loss term keeps the model’s general capabilities intact by encouraging its safe outputs to remain aligned with the original distribution.

### v3

@@ 1 Introduction
Figure 1: RepBend bends the safe and unsafe representations spaces of the model to be far and distinguishable from each other (Right). Before applying RepBend, the model cannot tell if “To make a bomb …” is safe or unsafe, as those regions in the activation space are close and intertwined to each other (Left).
@@ 3 Representation Bending
The core idea of representation bending is to “bend” the representation space of the model, ensuring that safe and unsafe states – latent knowledge encoded in the activation space – become distinctly separated and distant from each other. Specifically, representation bending pushes away the representations associated with harmful states from representations linked to safe behaviors, enhancing the model’s ability to clearly differentiate between safe and unsafe outcomes (Figure 1).
@@
RepBend is shown in Algorithm 1. RepBend tries to change the representations of model $M$ by minimizing a loss function over a fixed number of steps during fine-tuning. For improved performance, RepBend updates the model parameters with LoRA Hu et al. (2021). Model with LoRA is referred to as (and can be seen as an independent) model $M^{\prime}$. Since we only update the LoRA parameters, $M^{\prime}$ at the end of the algorithm would be the desired safe model.
@@
Line 12 shows four terms in the loss function. First loss term keeps the model’s representation close to the safe representations (small L2 difference); this can be seen as “retain loss.” Second loss term is the opposite of the first one: it pushes the representations of the model to be far away from unsafe representations (large L2 difference). This term can be seen as “forget loss.” These first two loss terms together separate the unsafe representation space from the safe representation space. Third, the cosine similarity loss term stabilizes the model responses under the unsafe queries toward a single “refusal-like” direction (e.g., “I am sorry I cannot help” or “As an AI assistant, I am unable to …”), helping the model consistently reject them rather than producing random or inconsistent outputs. The cosine similarity loss stabilizes the representation bending process by encouraging harmful representations to align closely with one another. Finally, the KL divergence loss term keeps the model’s general capabilities intact by encouraging its safe outputs to remain aligned with the original distribution.